OpenStudy (anonymous):

radioactive decay question?? The rate of decay of a radioactive substance is proportional to the amount present. Four years ago there were 12 grams of substance. Now there are 8 grams. How many grams will there be 8 years from now?

OpenStudy (anonymous):

Please help!

OpenStudy (kropot72):

The amount four years ago is determined by the equation \[N _{t}=N _{0}e ^{-kt}=12\ .............(1)\] The amount now is determined by the equation \[N _{t+4}=N _{0}e ^{-k(t+4)}=8\ .........(2)\] Dividing equation (1) by equation (2) gives \[\frac{e ^{-kt}}{e ^{-k(t+4)}}=\frac{12}{8}\] from which we get \[e ^{4k}=1.5\] When you solve for the decay constant k, you can set up another pair of equations for the weight now and the weight 8 years from now and solve for the weight in 8 years.